# FlashSpeech: Efficient Zero-Shot Speech Synthesis

### v1

@@ 1 Introduction
In recent years, the landscape of speech synthesis has been transformed by the advent of large-scale generative models. Consequently, the latest research efforts have achieved notable advancements in zero-shot speech synthesis systems by significantly increasing the size of both datasets and models. Zero-shot speech synthesis, such as text-to-speech (TTS), voice conversion (VC) and Editing, aims to generate speech that incorporates unseen speaker characteristics from a reference audio segment during inference, without the need for additional training. Current advanced zero-shot speech synthesis systems typically leverage language models (LMs) Wang et al. (2023a); Yang et al. (2023); Zhang et al. (2023); Kharitonov et al. (2023); Wang et al. (2023b); Peng et al. (2024); Kim et al. (2024) and diffusion-style models Shen et al. ([n. d.]); Kim et al. (2023b); Le et al. (2024); Jiang et al. (2023b) for in-context speech generation on the large-scale dataset. However, the generation process of these methods needs a long-time iteration. For example, VALL-E Wang et al. (2023a) builds on the language model to predict 75 audio token sequences for a 1-second speech, in its first-stage autoregressive (AR) token sequence generation. When using a non-autoregressive (NAR) latent diffusion model Rombach et al. (2022) based framework, NaturalSpeech 2 Shen et al. ([n. d.]) still requires 150 sampling steps. As a result, although these methods can produce human-like speech, they require significant computational time and cost. Some efforts have been made to accelerate the generation process. Voicebox Le et al. (2024) adopts flow-matching Lipman et al. (2022) so that fewer sampling steps (NFE***NFE: number of function evaluations.: 64) can be achieved because of the optimal transport path. ClaM-TTS Kim et al. (2024) proposes a mel-codec with a superior compression rate and a latent language model that generates a stack of tokens at once. Although the slow generation speed issue has been somewhat alleviated, the inference speed is still far from satisfactory for practical applications. Moreover, the substantial computational time of these approaches leads to significant computational cost overheads, presenting another challenge.
@@
The fundamental limitation of speech generation stems from the intrinsic mechanisms of language models and diffusion models, which require considerable time either auto-regressively or through a large number of denoising steps. Hence, the primary objective of this work is to accelerate inference speed and reduce computational costs while preserving generation quality at levels comparable to the prior research. In this paper, we propose FlashSpeech as the next step towards efficient zero-shot speech synthesis. To address the challenge of slow generation speed, we leverage the latent consistency model (LCM) Luo et al. (2023), a recent advancement in generative models. Building upon the previous non-autoregressive TTS system Shen et al. ([n. d.]), we adopt the encoder of a neural audio codec to convert speech waveforms into latent vectors as the training target for our LCM. To train this model, we propose a novel technique called adversarial consistency training, which utilizes the capabilities of pre-trained speech language models Chen et al. (2022b); Hsu et al. (2021); Baevski et al. (2020) as discriminators. This facilitates the transfer of knowledge from large pre-trained speech language models to speech generation tasks, efficiently integrating adversarial and consistency training to improve performance. The LCM is conditioned on prior vectors obtained from a phoneme encoder, a prompt encoder, and a prosody generator. Furthermore, we demonstrate that our proposed prosody generator leads to more diverse expressions and prosody while preserving stability.
@@ 2.1 Large-Scale Speech Synthesis
Motivated by the success of the large language model, the speech research community has recently shown increasing interest in scaling the sizes of model and training data to bolster generalization capabilities, producing natural speech with diverse speaker identities and prosody under zero-shot settings. The pioneering work is VALL-E Wang et al. (2023a), which adopts the Encodec Défossez et al. (2022) to discretize the audio waveform into tokens. Therefore, a language model can be trained via in-context learning that can generate the target utterance where the style is consistent with prompt utterance. However, generating audio in such an autoregressive manner Wang et al. (2023b); Peng et al. (2024)can lead to unstable prosody, word skipping, and repeating issues Ren et al. (2020); Tan et al. (2021); Shen et al. ([n. d.]). To ensure the robustness of the system, non-autoregressive methods such as NaturalSpeech2 Shen et al. ([n. d.]) and Voicebox Le et al. (2024) utilize diffusion-style model (VP-diffusion Song et al. (2020) or flow-matching Lipman et al. (2022)) to learn the distribution of a continuous intermediate vector such as mel-spectrogram or latent vector of codec. Both LM-based methods Zhao et al. (2023) and diffusion-based methods show superior performance in speech generation tasks. However, their generation is slow due to the iterative computation. Considering that many speech generation scenarios require real-time inference and low computational costs, we employ the latent consistency model for large-scale speech generation that inference with one or two steps while maintaining high audio quality.
@@ 3.4.1 Analysis of Prosody Prediction
Previous regression methods for prosody prediction Ren et al. (2020); Shen et al. ([n. d.]), due to their deterministic mappings and assumptions of unimodal distribution, often fail to capture the inherent diversity and expressiveness of human speech prosody. This leads to predictions that lack variation and can appear over-smoothed. On the other hand, diffusion methods Le et al. (2024); Li et al. (2023) for prosody prediction offer a promising alternative by providing greater prosody diversity. However, they come with challenges regarding stability, and the potential for unnatural prosody. Additionally, the iterative inference process in DMs requires a significant number of sampling steps that may also hinder real-time application. Meanwhile, LM-based methods Jiang et al. (2024a); Wang et al. (2023a) also need a long time for inference. To alleviate these issues, our prosody generator consists of a prosody regression module and a prosody refinement module to enhance the diversity of prosody regression results with efficient one-step consistency model sampling.
@@ 3.5.2 Voice Conversion
Voice conversion aims to convert the source audio into the target audio using the speaker’s voice of the reference audio. Following Shen et al. ([n. d.]); Preechakul et al. (2022), we first apply the reverse of ODE to diffuse the source audio into a starting point that still maintains some information in the source audio. After that, we run the sampling process from this starting point with the reference audio as $\mathbf{z}_{prompt}$ and condition $c$. The condition $c$ uses the phoneme and duration from the source audio and the pitch is predicted by the prosody generator. This method allows for zero-shot voice conversion while preserving the linguistic content of the source audio, and achieving the same timbre as the reference audio.
@@ 4.1.1 Data and Preprocessing
We use the English subset of Multilingual LibriSpeech (MLS) Pratap et al. (2020), including 44.5k hours of transcribed audiobook data and it contains 5490 distinct speakers. The audio data is resampled at a frequency of 16kHz. The input text is transformed into a sequence of phonemes through grapheme-to-phoneme conversion Sun et al. (2019) and then we use our internal alignment tool aligned with speech to obtain the phoneme-level duration. We adopt a hop size of 200 for all frame-level features. The pitch sequence is extracted using PyWorld******https://github.com/JeremyCCHsu/Python-Wrapper-for-World-Vocoder. we adopt Encodec Défossez et al. (2022) as our audio codec. We use a modified version *********https://github.com/yangdongchao/UniAudio/tree/main/codec and train it on MLS. We use the dense features extracted before the residual quantization layer as our latent vector $z$.
@@ 4.1.3 Model Details
The model structures of the prompt encoder and phoneme encoder are followShen et al. ([n. d.]). The neural function part in LCM is almost the same as the Shen et al. ([n. d.]). We rescale the sinusoidal position embedding in the neural function part by a factor of 1000. As for the prosody generator, we adopt 30 non-casual wavenet Oord et al. (2016) layers for the neural function part in the prosody refinement module and the same configurations for prosody regression parts in Shen et al. ([n. d.]). And we set $\alpha=0.2$ for the prosody refinement module empirically. For the discriminator’s head, we stack 5 convolutional layers with weight normalization Salimans and Kingma (2016) for binary classification.
@@ 4.2 Evaluation Metrics
  • RTF: Real-time-factor (RTF) measures the time taken for the system to generate one second of speech. This metric is crucial for evaluating the efficiency of our system, particularly for applications requiring real-time processing. We measure the time of our system end-to-end on an NVIDIA V100 GPU following Shen et al. ([n. d.]).
  • Sim-O and Sim-R: These metrics assess the speaker similarity. Sim-R measures the objective similarity between the synthesized speech and the reconstruction reference speech through the audio codec, using features embedding extracted from the pre-trained speaker verification model Wang et al. (2023a); Kim et al. (2024)https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification. Sim-O is calculated with the original reference speech. Higher scores in Sim-O and Sim-R indicate a higher speaker similarity.
  • WER (Word Error Rate): To evaluate the accuracy and clarity of synthesized speech from the TTS system, we employ the Automatic Speech Recognition (ASR) model Wang et al. (2023a) https://huggingface.co/facebook/hubert-large-ls960-ft to transcribe generated audio. The discrepancies between these transcriptions and original texts are quantified using the Word Error Rate (WER), a crucial metric indicating intelligibility and robustness.
  • CMOS, SMOS, UTMOS: we rank the comparative mean option score (CMOS) and similarity mean option score (SMOS) using mturk. The prompt for CMOS refers to ’Please focus on the audio quality and naturalness and ignore other factors.’. The prompt for SMOS refers to ’Please focus on the similarity of the speaker to the reference, and ignore the differences of content, grammar or audio quality.’ Each audio has been listened to by at least 10 listeners. UTMOS Saeki et al. (2022) is a Speech MOS predictorhttps://github.com/tarepan/SpeechMOS to measure the naturalness of speech. We use it in ablation studies which reduced the cost for evaluation.
@@ 4.3.1 Evaluation Baselines
  • VALL-E Wang et al. (2023a): VALL-E predicts codec tokens using both AR and NAR models. RTFIn CLaM-TTS and Voicebox, they report the inference time for generating 10 seconds of speech. Therefore, we divide by 10 to obtain the time for generating 1 second of speech (RTF). is obtained from Kim et al. (2024); Le et al. (2024). We use our reproduced results for MOS, Sim, and WER. Additionally, we do a preference test with their official demo.
@@
  • NaturalSpeech2 Shen et al. ([n. d.]): NaturalSpeech2 uses a latent diffusion model to predict latent features of codec. The RTF is from the original paper. the Sim, WER and samples for MOS are obtained through communication with the authors. We also do a preference test with their official demo.
  • Mega-TTS Jiang et al. (2023a)Since we do not find any audio samples for Mega-TTS2 Jiang et al. (2024b) under the 3-second cross-sentence setting, we are not able to compare with them.: Mega-TTS uses both language model and GAN to predict mel-spectrogram. We obtain RTF from mobilespeech Ji et al. (2024) and WER from the original paper. We do a preference test with their official demo.
@@ 4.4.1 Ablation studies of LCM
We explored the impact of different pre-trained models in adversarial training on UTMOS and Sim-O. As shown in the table 2, the baseline, which employs consistency training alone, achieved a UTMOS of 3.62 and a Sim-O of 0.45. Incorporating adversarial training using wav2vec2-largehttps://huggingface.co/facebook/wav2vec2-large, hubert-largehttps://huggingface.co/facebook/hubert-large-ll60k, and wavlm-largehttps://huggingface.co/microsoft/wavlm-large as discriminators significantly improved both UTMOS and Sim-O scores. Notably, the application of adversarial training with Wavlm-large achieved the highest scores (UTMOS: 4.00, Sim-O: 0.52), underscoring the efficacy of this pre-trained model in enhancing the quality and speaker similarity of synthesized speech. Additionally, without using the audio prompt’s feature as a condition the discriminator shows a slight decrease in performance (UTMOS: 3.97, Sim-O: 0.51), highlighting the importance of conditional features in guiding the adversarial training process.
@@ 4.5 Evaluation Results for Voice Conversion
In this section, we present the evaluation results of our voice conversion system, FlashSpeech, in comparison with state-of-the-art methods, including YourTTS https://github.com/coqui-ai/TTS Casanova et al. (2022) and DDDM-VC https://github.com/hayeong0/DDDM-VC Choi et al. (2024). We conduct the experiments with their official checkpoints in our internal test set.

### v3

@@ 1 Introduction
In recent years, the landscape of speech synthesis has been transformed by the advent of large-scale generative models. Consequently, the latest research efforts have achieved notable advancements in zero-shot speech synthesis systems by significantly increasing the size of both datasets and models. Zero-shot speech synthesis, such as text-to-speech (TTS), voice conversion (VC) and Editing, aims to generate speech that incorporates unseen speaker characteristics from a reference audio segment during inference, without the need for additional training. Current advanced zero-shot speech synthesis systems typically leverage language models (LMs) Wang et al. (2023a); Yang et al. (2023); Zhang et al. (2023); Kharitonov et al. (2023); Wang et al. (2023b); Peng et al. (2024); Kim et al. (2024) and diffusion-style models Shen et al. (2024); Kim et al. (2023b); Le et al. (2023); Jiang et al. (2023b) for in-context speech generation on the large-scale dataset. However, the generation process of these methods needs a long-time iteration. For example, VALL-E Wang et al. (2023a) builds on the language model to predict 75 audio token sequences for a 1-second speech, in its first-stage autoregressive (AR) token sequence generation. When using a non-autoregressive (NAR) latent diffusion model Rombach et al. (2022) based framework, NaturalSpeech 2 Shen et al. (2024) still requires 150 sampling steps. As a result, although these methods can produce human-like speech, they require significant computational time and cost. Some efforts have been made to accelerate the generation process. Voicebox Le et al. (2023) adopts flow-matching Lipman et al. (2022) so that fewer sampling steps (NFE [^1] NFE: number of function evaluations.: 64) can be achieved because of the optimal transport path. ClaM-TTS Kim et al. (2024) proposes a mel-codec with a superior compression rate and a latent language model that generates a stack of tokens at once. Although the slow generation speed issue has been somewhat alleviated, the inference speed is still far from satisfactory for practical applications. Moreover, the substantial computational time of these approaches leads to significant computational cost overheads, presenting another challenge.
@@
The fundamental limitation of speech generation stems from the intrinsic mechanisms of language models and diffusion models, which require considerable time either auto-regressively or through a large number of denoising steps. Hence, the primary objective of this work is to accelerate inference speed and reduce computational costs while preserving generation quality at levels comparable to the prior research. In this paper, we propose FlashSpeech as the next step towards efficient zero-shot speech synthesis. To address the challenge of slow generation speed, we leverage the latent consistency model (LCM) Luo et al. (2023), a recent advancement in generative models. Building upon the previous non-autoregressive TTS system Shen et al. (2024), we adopt the encoder of a neural audio codec to convert speech waveforms into latent vectors as the training target for our LCM. To train this model, we propose a novel technique called adversarial consistency training, which utilizes the capabilities of pre-trained speech language models Chen et al. (2022b); Hsu et al. (2021); Baevski et al. (2020) as discriminators. This facilitates the transfer of knowledge from large pre-trained speech language models to speech generation tasks, efficiently integrating adversarial and consistency training to improve performance. The LCM is conditioned on prior vectors obtained from a phoneme encoder, a prompt encoder, and a prosody generator. Furthermore, we demonstrate that our proposed prosody generator leads to more diverse expressions and prosody while preserving stability.
@@ 2.1 Large-Scale Speech Synthesis
Motivated by the success of the large language model, the speech research community has recently shown increasing interest in scaling the sizes of model and training data to bolster generalization capabilities, producing natural speech with diverse speaker identities and prosody under zero-shot settings. The pioneering work is VALL-E Wang et al. (2023a), which adopts the Encodec Défossez et al. (2022) to discretize the audio waveform into tokens. Therefore, a language model can be trained via in-context learning that can generate the target utterance where the style is consistent with prompt utterance. However, generating audio in such an autoregressive manner Wang et al. (2023b); Peng et al. (2024)can lead to unstable prosody, word skipping, and repeating issues Ren et al. (2020); Tan et al. (2021); Shen et al. (2024). To ensure the robustness of the system, non-autoregressive methods such as NaturalSpeech2 Shen et al. (2024) and Voicebox Le et al. (2023) utilize diffusion-style model (VP-diffusion Song et al. (2020) or flow-matching Lipman et al. (2022)) to learn the distribution of a continuous intermediate vector such as mel-spectrogram or latent vector of codec. Both LM-based methods Zhao et al. (2023) and diffusion-based methods show superior performance in speech generation tasks. However, their generation is slow due to the iterative computation. Considering that many speech generation scenarios require real-time inference and low computational costs, we employ the latent consistency model for large-scale speech generation that inference with one or two steps while maintaining high audio quality.
@@ 3.4.1 Analysis of Prosody Prediction
Previous regression methods for prosody prediction Ren et al. (2020); Shen et al. (2024), due to their deterministic mappings and assumptions of unimodal distribution, often fail to capture the inherent diversity and expressiveness of human speech prosody. This leads to predictions that lack variation and can appear over-smoothed. On the other hand, diffusion methods Le et al. (2023); Li et al. (2023) for prosody prediction offer a promising alternative by providing greater prosody diversity. However, they come with challenges regarding stability, and the potential for unnatural prosody. Additionally, the iterative inference process in DMs requires a significant number of sampling steps that may also hinder real-time application. Meanwhile, LM-based methods Jiang et al. (2024a); Wang et al. (2023a) also need a long time for inference. To alleviate these issues, our prosody generator consists of a prosody regression module and a prosody refinement module to enhance the diversity of prosody regression results with efficient one-step consistency model sampling.
@@ 3.5.2 Voice Conversion
Voice conversion aims to convert the source audio into the target audio using the speaker’s voice of the reference audio. Following Shen et al. (2024); Preechakul et al. (2022), we first apply the reverse of ODE to diffuse the source audio into a starting point that still maintains some information in the source audio. After that, we run the sampling process from this starting point with the reference audio as $\mathbf{z}_{prompt}$ and condition $c$. The condition $c$ uses the phoneme and duration from the source audio and the pitch is predicted by the prosody generator. This method allows for zero-shot voice conversion while preserving the linguistic content of the source audio, and achieving the same timbre as the reference audio.
@@ 4.1.1 Data and Preprocessing
We use the English subset of Multilingual LibriSpeech (MLS) Pratap et al. (2020), including 44.5k hours of transcribed audiobook data and it contains 5490 distinct speakers. The audio data is resampled at a frequency of 16kHz. The input text is transformed into a sequence of phonemes through grapheme-to-phoneme conversion Sun et al. (2019) and then we use our internal alignment tool aligned with speech to obtain the phoneme-level duration. We adopt a hop size of 200 for all frame-level features. The pitch sequence is extracted using PyWorld [^2] https://github.com/JeremyCCHsu/Python-Wrapper-for-World-Vocoder. we adopt Encodec Défossez et al. (2022) as our audio codec. We use a modified version  [^3] https://github.com/yangdongchao/UniAudio/tree/main/codec and train it on MLS. We use the dense features extracted before the residual quantization layer as our latent vector $z$.
@@ 4.1.3 Model Details
The model structures of the prompt encoder and phoneme encoder are followShen et al. (2024). The neural function part in LCM is almost the same as the Shen et al. (2024). We rescale the sinusoidal position embedding in the neural function part by a factor of 1000. As for the prosody generator, we adopt 30 non-casual wavenet Oord et al. (2016) layers for the neural function part in the prosody refinement module and the same configurations for prosody regression parts in Shen et al. (2024). And we set $\alpha=0.2$ for the prosody refinement module empirically. For the discriminator’s head, we stack 5 convolutional layers with weight normalization Salimans and Kingma (2016) for binary classification.
@@ 4.2 Evaluation Metrics
  • RTF: Real-time-factor (RTF) measures the time taken for the system to generate one second of speech. This metric is crucial for evaluating the efficiency of our system, particularly for applications requiring real-time processing. We measure the time of our system end-to-end on an NVIDIA V100 GPU following Shen et al. (2024).
  • Sim-O and Sim-R: These metrics assess the speaker similarity. Sim-R measures the objective similarity between the synthesized speech and the reconstruction reference speech through the audio codec, using features embedding extracted from the pre-trained speaker verification model Wang et al. (2023a); Kim et al. (2024) [^4] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification. Sim-O is calculated with the original reference speech. Higher scores in Sim-O and Sim-R indicate a higher speaker similarity.
  • WER (Word Error Rate): To evaluate the accuracy and clarity of synthesized speech from the TTS system, we employ the Automatic Speech Recognition (ASR) model Wang et al. (2023a)  [^5] https://huggingface.co/facebook/hubert-large-ls960-ft to transcribe generated audio. The discrepancies between these transcriptions and original texts are quantified using the Word Error Rate (WER), a crucial metric indicating intelligibility and robustness.
  • CMOS, SMOS, UTMOS: we rank the comparative mean option score (CMOS) and similarity mean option score (SMOS) using mturk. The prompt for CMOS refers to ’Please focus on the audio quality and naturalness and ignore other factors.’. The prompt for SMOS refers to ’Please focus on the similarity of the speaker to the reference, and ignore the differences of content, grammar or audio quality.’ Each audio has been listened to by at least 10 listeners. UTMOS Saeki et al. (2022) is a Speech MOS predictor [^6] https://github.com/tarepan/SpeechMOS to measure the naturalness of speech. We use it in ablation studies which reduced the cost for evaluation.
@@ 4.3.1 Evaluation Baselines
  • VALL-E Wang et al. (2023a): VALL-E predicts codec tokens using both AR and NAR models. RTF [^7] In CLaM-TTS and Voicebox, they report the inference time for generating 10 seconds of speech. Therefore, we divide by 10 to obtain the time for generating 1 second of speech (RTF). is obtained from Kim et al. (2024); Le et al. (2023). We use our reproduced results for MOS, Sim, and WER. Additionally, we do a preference test with their official demo.
@@
  • NaturalSpeech2 Shen et al. (2024): NaturalSpeech2 uses a latent diffusion model to predict latent features of codec. The RTF is from the original paper. the Sim, WER and samples for MOS are obtained through communication with the authors. We also do a preference test with their official demo.
  • Mega-TTS Jiang et al. (2023a) [^8] Since we do not find any audio samples for Mega-TTS2 Jiang et al. (2024b) under the 3-second cross-sentence setting, we are not able to compare with them.: Mega-TTS uses both language model and GAN to predict mel-spectrogram. We obtain RTF from mobilespeech Ji et al. (2024) and WER from the original paper. We do a preference test with their official demo.
@@ 4.4.1 Ablation studies of LCM
We explored the impact of different pre-trained models in adversarial training on UTMOS and Sim-O. As shown in the table 2, the baseline, which employs consistency training alone, achieved a UTMOS of 3.62 and a Sim-O of 0.45. Incorporating adversarial training using wav2vec2-large [^9] https://huggingface.co/facebook/wav2vec2-large, hubert-large [^10] https://huggingface.co/facebook/hubert-large-ll60k, and wavlm-large [^11] https://huggingface.co/microsoft/wavlm-large as discriminators significantly improved both UTMOS and Sim-O scores. Notably, the application of adversarial training with Wavlm-large achieved the highest scores (UTMOS: 4.00, Sim-O: 0.52), underscoring the efficacy of this pre-trained model in enhancing the quality and speaker similarity of synthesized speech. Additionally, without using the audio prompt’s feature as a condition the discriminator shows a slight decrease in performance (UTMOS: 3.97, Sim-O: 0.51), highlighting the importance of conditional features in guiding the adversarial training process.
@@ 4.5 Evaluation Results for Voice Conversion
In this section, we present the evaluation results of our voice conversion system, FlashSpeech, in comparison with state-of-the-art methods, including YourTTS  [^12] https://github.com/coqui-ai/TTS Casanova et al. (2022) and DDDM-VC  [^13] https://github.com/hayeong0/DDDM-VC Choi et al. (2024). We conduct the experiments with their official checkpoints in our internal test set.